# Flagged Extensions and Numerical Simulations for Quantum Channel Capacity: Bridging Theory and Computation

### v1

@@ 2 Mathematical Formulation
These theorems and proposition provide a toolbox for analyzing channel capacities. Theorem 2.1 is established theory, while Theorem 2.2 and Proposition 2.3 represent newer insights and will be applied in our analysis. In the following, we validate these results and explore their consequences via numerical simulations.
@@ 3 Numerical Results
In this section, we present our numerical simulations of coherent information for three prominent quantum channels: (1) the amplitude damping channel, (2) the depolarizing channel, and (3) the flagged depolarizing channel. Our goal is to investigate how the maximum achievable one-shot coherent information changes with the channel noise parameter, validating the theoretical discussion in previous sections.
@@ 3.1 Amplitude Damping Channel
We begin with the amplitude damping channel, parameterized by the damping probability $\gamma\in[0,1]$. In the limit $\gamma\approx 0$, the channel behaves nearly like an identity map (i.e., minimal noise), while larger values of $\gamma$ signify stronger damping of the excited state $\lvert 1\rangle$ to $\lvert 0\rangle$. Figure 1 shows our simulation for $\gamma$ ranging from 00 to $0.6$.
As $\gamma$ increases from 00, the coherent information decreases, reflecting the growing impact of damping. At very low $\gamma$, the channel is close to ideal, and one can achieve coherent information values exceeding three bits for single-letter uses (depending on the input pure state). However, as the damping probability rises, the coherent information eventually stabilizes and fluctuates around smaller values, indicating that the environment is capturing a significant portion of the information. The amplitude damping channel is known to be degradable for all $\gamma\leq 1$, implying its capacity is single-letter in principle; our numerical results capture the progressive decline in achievable one-shot coherent information due to stronger damping.
@@ 3.2 Depolarizing Channel
Next, we examine the depolarizing channel, where each qubit is replaced by the maximally mixed state $\tfrac{I}{2}$ with probability $p$. Figure 2 illustrates the maximum single-letter coherent information as a function of $p\in[0,0.6]$.
Because pure-state inputs to the depolarizing channel often yield a negative one-shot coherent information (the environment can learn more about the state than the receiver), our plot indeed shows that coherent information takes negative or near-zero values for even moderate $p$. This aligns with theoretical expectations that the single-letter coherent information of a qubit depolarizing channel is typically below zero for any $p>0$. Hence, while the channel may still have a positive capacity when considering multi-letter or entangled inputs across multiple uses, the single-letter scenario we simulate here remains dominated by noise, resulting in nonpositive values.
@@ 3.3 Flagged Depolarizing Channel
Finally, we investigate the flagged depolarizing channel, which augments the original map with an ancillary “flag” system to indicate which noise event occurred. Such an extension is intended to yield higher (or at least no smaller) coherent information compared to the plain depolarizing channel, because the receiver gains additional classical information about the noise branch. Figure 3 plots the maximum single-letter coherent information for $p$ up to $0.6$.
@@
Discussion.
These numerical results highlight that while single-letter coherent information can be positive for low-noise channels like amplitude damping at small $\gamma$, it drops quickly once the noise level increases. For the depolarizing and flagged depolarizing channels, single-letter coherent information remains near zero or negative, consistent with known results that depolarizing noise is extremely detrimental for pure-state inputs. In a multi-letter or entanglement-assisted regime, one might recover a larger effective capacity, but the single-letter scenario is generally pessimistic about quantum transmission in these noisy settings. Overall, our simulations confirm the theoretical predictions and provide a practical window into the channel capacity limitations discussed in Sections 2.1 and 2.2.
@@ 4 Conclusions
In this paper, we investigated the capacities of noisy quantum channels through a combined analytical and numerical approach. We introduced novel flagged extension techniques that embed a channel into a higher-dimensional space, thereby yielding single-letter upper bounds on the quantum and private capacities. Our theoretical results, which include the Quantum Capacity Theorem (Theorem 2.1), the Flagged Extension Upper Bound (Theorem 2.2), and the Zero-Capacity Noise Threshold (Proposition 2.3), provide rigorous tools for understanding the limitations imposed by noise.
Furthermore, our numerical simulations—conducted for the amplitude damping, depolarizing, and flagged depolarizing channels—validate these theoretical bounds. The simulations highlight that while the depolarizing channel exhibits near-zero or negative single-letter coherent information, the amplitude damping channel (being degradable) presents a positive albeit modest coherent information in low-noise regimes. These results underscore the challenges in extracting quantum capacity from pure-state inputs and indicate that further investigation using multi-letter or mixed-state strategies could be worthwhile.
Future work may extend the current approach by incorporating mixed-state inputs, alternative variational ansätze, or experimental validation on real quantum hardware. Such extensions could further narrow the gap between our theoretical bounds and the true operational capacities of noisy quantum channels.

### v2

@@ 2 Mathematical Formulation
These theorems and proposition provide a toolbox for analyzing channel capacities. Theorem 2.4 and Proposition 2.5 represent newer insights and will be applied in our analysis. In the following, I validate these results and explore their consequences via numerical simulations.
3 Methods and Reproducibility
I study three canonical qubit noise models: amplitude damping (AD) with damping $\gamma\in[0,1]$; generalized amplitude damping (GADC) with parameters $(\gamma,N_{\mathrm{th}})$ describing a thermal bath with population $N_{\mathrm{th}}\in[0,1]$; and depolarizing (Dep) with probability $p\in[0,1]$. The GADC Kraus form, degradability/anti-degradability regimes, and entanglement-breaking thresholds follow [6]. Flagged extensions used for capacity upper bounds follow the “quantum flags” framework of [3, 7]. Standard capacity notions and notation follow [21, 2].
For a channel $\mathcal{N}$ we compute (i) the one-shot coherent information $I_{\mathrm{c}}(\rho,\mathcal{N})=H(\mathcal{N}(\rho))-H(\mathcal{N}^{c}(\rho))$ maximized over inputs $\rho$; (ii) the reverse coherent information $I_{\mathrm{rc}}(\rho,\mathcal{N})$ reported as a comparative indicator on AD/GADC; (iii) the entanglement-assisted mutual information $I(A;B)$ for scale; and (iv) upper bounds from flagged extensions for convex mixtures [3, 7] and from the additive SDP bound $Q_{\Gamma}$ [18].
Inputs are restricted to pure qubit states $|\psi(\theta,\phi)\rangle=\cos(\tfrac{\theta}{2})|0\rangle+e^{i\phi}\sin(\tfrac{\theta}{2})|1\rangle$ with $(\theta,\phi)\in[0,\pi]\times[0,2\pi)$. We use a two-stage scheme: (1) global sampling over $(\theta,\phi)$ via Sobol/Latin-hypercube to select candidates; (2) local derivative-free refinement (Powell/Nelder–Mead) from the top seeds. Entropies are evaluated from eigenvalues with $10^{-12}$ clipping to avoid numerical artifacts; termination requires successive objective change $10^{-9}$. For AD, optima align with the energy basis; for GADC, off-axis optima can occur at intermediate $(\gamma,N_{\mathrm{th}})$, consistent with [6].
AD: $\gamma\in[0,0.9]$ with finer sampling near $\gamma\!\approx\!1/2$ to resolve the (anti-)degradability transition. GADC: $\gamma\in[0,0.9]$ for $N_{\mathrm{th}}\in\{0,0.1,0.2\}$ with regime annotations from [6]. Depolarizing: $p\in[0,0.35]$; we compare the hashing lower bound with flagged-extension upper bounds from [3, 7] and the SDP $Q_{\Gamma}$ [18].
For mixtures $\mathcal{N}=p\,\mathcal{U}+(1-p)\,\mathcal{M}$, we construct the flagged extension and evaluate its single-letter capacity as an explicit upper bound on $Q(\mathcal{N})$ (and $P(\mathcal{N})$). We reuse the degradability conditions and tight depolarizing-channel bounds from [3, 7].
The additive semidefinite-programming bound $Q_{\Gamma}$ is computed on the Choi operator $J(\mathcal{N})$; we solve the primal to tolerance $10^{-8}$ with a standard conic solver (CVX/SDP). Additivity ensures consistency across blocklengths and facilitates comparison with flagged-extension bounds [18].
Algorithm 1 Maximizing $I_{\mathrm{c}}$ over qubit inputs for a fixed channel $\mathcal{N}$
1:Input: channel $\mathcal{N}$ (Kraus ops or Choi), grid size $N_{\mathrm{grid}}$, seeds $N_{\mathrm{seeds}}$
2: Sample $(\theta_{i},\phi_{i})_{i=1}^{N_{\mathrm{grid}}}$ via Sobol/LH in $0,\pi]\times[0,2\pi)$
3:for $i=1$ to $N_{\mathrm{grid}}$ do
4:  $\rho_{i}\leftarrow|\psi(\theta_{i},\phi_{i})\rangle\!\langle\psi(\theta_{i},\phi_{i})|$
5:  Evaluate $I_{\mathrm{c}}(\rho_{i},\mathcal{N})$ using eigenvalue entropy with clipping
6:end for
7: Select top $N_{\mathrm{seeds}}$ points; for each, run local trust-region search on $(\theta,\phi)$
8: Return $\max$ over all refined values and the corresponding $\rho^{\star}$
@@ 4 Numerical Results
Table 1: Qubit depolarizing channel $D_{p}(\rho)=(1-p)\rho+p\,I/2$: comparison of lower and upper bounds (bits/use). Baseline lower bound is the hashing/coherent-information rate $1-H_{2}(p)-p\log_{2}3$. “FKG PRL (Eq. (11))” is the Fanizza–Kianvash–Giovannetti flagged-extension bound [3]; “Convex hull (Eq. (16))” is the convex hull of degradable-extension bounds [8]; “SDP / max-Rains” is the strong-converse bound equal to the max-Rains information [18, 19].
$p$ Hashing lower bound FKG PRL (Eq. (11)) Convex hull (Eq. (16)) SDP / max-Rains
0.05 0.6343549 0.757 0.741 0.812
0.10 0.3725082 0.607 0.592 0.670
0.15 0.1524153 0.462 0.450 0.541
Comparison.
At small noise (p=0.05p{=}0.05), the FKG flagged-extension bound improves substantially over earlier symmetric-side-channel and approximate-degradability converses, while its convex-hull refinement reduces the value by a few hundredths of a bit [3, 8]. Across $p\in\{0.05,0.10,0.15\}$, the convex-hull envelope remains slightly below the FKG curve, and the SDP/max-Rains bound sits above both in the low-noise regime [18, 19]. Relative to the hashing lower bound, the feasible interval (Upper - Lower) shrinks to roughly 0.110.110.390.39 bits in this range.
In this section, I present our numerical simulations of coherent information for three prominent quantum channels: (1) the amplitude damping channel, (2) the depolarizing channel, and (3) the flagged depolarizing channel. Our goal is to investigate how the maximum achievable one-shot coherent information changes with the channel noise parameter, validating the theoretical discussion in previous sections.
@@ 4.1 Amplitude Damping Channel
I begin with the amplitude damping channel, parameterized by the damping probability $\gamma\in[0,1]$. In the limit $\gamma\approx 0$, the channel behaves nearly like an identity map (i.e., minimal noise), while larger values of $\gamma$ signify stronger damping of the excited state $\lvert 1\rangle$ to $\lvert 0\rangle$. Figure 1 shows our simulation for $\gamma$ ranging from 0 to $0.6$.
As $\gamma$ increases from 0, the coherent information decreases, reflecting the growing impact of damping. At very low $\gamma$, the channel is close to ideal, and the maximized single-letter coherent information approaches $1$ bit per channel use (the qubit limit), depending on the input state. However, as the damping probability rises, the coherent information eventually stabilizes and fluctuates around smaller values, indicating that the environment is capturing a significant portion of the information. The amplitude-damping channel is degradable for $\gamma\leq\tfrac{1}{2}$ and anti-degradable for $\gamma\geq\tfrac{1}{2}$; in the degradable regime its quantum capacity equals its coherent information, while beyond the threshold the capacity vanishes. Our numerical results capture the progressive decline in achievable one-shot coherent information as $\gamma$ increases.
@@ 4.2 Depolarizing Channel
Next, I examine the depolarizing channel, where each qubit is replaced by the maximally mixed state $\tfrac{I}{2}$ with probability $p$. Figure 2 illustrates the maximum single-letter coherent information as a function of $p\in[0,0.6]$.
Because pure-state inputs to the depolarizing channel often yield a negative one-shot coherent information (the environment can learn more about the state than the receiver), our plot indeed shows that coherent information takes negative or near-zero values for even moderate $p$. This aligns with theoretical expectations that the single-letter coherent information of a qubit depolarizing channel is typically below zero for any $p>0$. Hence, while the channel may still have a positive capacity when considering multi-letter or entangled inputs across multiple uses, the single-letter scenario I simulate here remains dominated by noise, resulting in nonpositive values.
@@ 4.3 Flagged Depolarizing Channel
Finally, I investigate the flagged depolarizing channel, which augments the original map with an ancillary “flag” system to indicate which noise event occurred. Such an extension is intended to yield higher (or at least no smaller) coherent information compared to the plain depolarizing channel, because the receiver gains additional classical information about the noise branch. Figure 3 plots the maximum single-letter coherent information for $p$ up to $0.6$.
@@
These numerical results highlight that while single-letter coherent information can be positive for low-noise channels like amplitude damping at small $\gamma$, it drops quickly once the noise level increases. For the depolarizing and flagged depolarizing channels, single-letter coherent information remains near zero or negative, consistent with known results that depolarizing noise is extremely detrimental for pure-state inputs. In a multi-letter or entanglement-assisted regime, one might recover a larger effective capacity, but the single-letter scenario is generally pessimistic about quantum transmission in these noisy settings. Overall, our simulations confirm the theoretical predictions and provide a practical window into the channel capacity limitations discussed in and 2.4.
@@ 5 Conclusions
I presented a unified analytical–numerical study of quantum and private communication over noisy qubit channels, with three main outcomes. First, on the theory side, I formalized a flagged–extension framework that yields computable, single-letter upper bounds whenever the non-unitary branch of the mixture is degradable. In this degradable-branch regime the flagged extension itself is degradable, so its quantum capacity equals its coherent information and upper bounds the capacity of the original channel. This clarifies the scope of flagged approaches and avoids over-claims in the fully general CPTP case. Our statements are consistent with and anchored in the standard single-letterization of degradable channels and the known degradability/anti-degradability threshold of the amplitude-damping (AD) family at $\gamma=\tfrac{1}{2}$ (degradable for $\gamma\leq\tfrac{1}{2}$, anti-degradable for $\gamma\geq\tfrac{1}{2}$).
Second, on the comparative numerics, I produced color/gray-safe capacity comparison curves for AD, generalized AD (GADC), and the qubit depolarizing channel. For AD (and low-bath GADC), the optimized coherent information matches the single-letter capacity in the degradable region and collapses as the anti-degradable threshold is crossed, in agreement with the structural results for degradable channels. For the depolarizing channel, I contrasted the hashing lower bound $I_{c}(\mathcal{D}_{p})=1-H_{2}(p)-p\log_{2}3$ with a semidefinite-programming (max-Rains) upper bound; the persistent gap in the intermediate-noise regime visualizes the standing uncertainty around the exact $Q(\mathcal{D}_{p})$ despite decades of progress.
Third, I complemented coherent-information plots with two additional, operationally meaningful benchmarks: (i) the reverse coherent information (RCI), which is an achievable lower bound based on reverse-reconciliation protocols, and (ii) an SDP (max-Rains) upper bound that is additive and efficiently computable. Together, these bounds bracket the attainable rates and turn the comparative figures into decision tools for parameter regimes of practical interest.
Acknowledgements
The author would like to thank David Mitchell for helpful editorial guidance and constructive comments that improved the clarity and presentation of the manuscript. The author also thanks the referees for their careful reading and insightful suggestions, which significantly strengthened both the theoretical discussion and the reproducibility of the numerical results.
Data availability statement
The data and code supporting the findings of this work are openly available at the following repository: https://github.com/vnourozi/quantum-channel-capacity-simulator.
Appendix A: Detailed proof ingredients for Theorem 2.4
Stinespring normal form.
Let $V_{U}$ and $V_{M}$ be Stinespring isometries for $\mathcal{U}$ and $\mathcal{M}$ acting as $V_{U}:\mathcal{H}_{A}\to\mathcal{H}_{B}\otimes\mathcal{H}_{E_{U}}$ and $V_{M}:\mathcal{H}_{A}\to\mathcal{H}_{B}\otimes\mathcal{H}_{E_{M}}$. Define $W=\sqrt{p}\,|0\rangle_{F}\otimes V_{U}\oplus\sqrt{1-p}\,|1\rangle_{F}\otimes V_{M}$. Then $\widetilde{\mathcal{N}}(\rho)=\mathrm{Tr}_{E}[W\rho W^{\dagger}]$ and $\widetilde{\mathcal{N}}^{c}(\rho)=\mathrm{Tr}_{B}[W\rho W^{\dagger}]$.
Explicit degrading map.
Let $\Delta$ be a degrading map for $\mathcal{M}$, i.e., $\mathcal{M}^{c}=\Delta\circ\mathcal{M}$. Set
$\mathsf{D}(\sigma_{FB})=|0\rangle\!\langle 0|_{F}\otimes\mathsf{D}_{0}(\langle 0|\sigma_{FB}|0\rangle_{F})+|1\rangle\!\langle 1|_{F}\otimes\mathsf{D}_{1}(\langle 1|\sigma_{FB}|1\rangle_{F}),$
with $\mathsf{D}_{0}(\cdot)=\mathcal{U}^{c}(\rho_{\star})$ (constant map) and $\mathsf{D}_{1}=\Delta$. Then $\mathsf{D}$ is CPTP and satisfies $\mathsf{D}\circ\widetilde{\mathcal{N}}=\widetilde{\mathcal{N}}^{c}$.
Choi positivity.
The Choi operator of $\mathsf{D}$ is block-diagonal in the flag, with a rank-1 block for $\mathsf{D}_{0}$ and the Choi operator of $\Delta$ for $\mathsf{D}_{1}$, hence positive and trace-preserving.
Specializations.
(i) For amplitude damping with $\gamma\leq 1/2$, a Kraus pair yields the known degrader (transpose channel); plugging it into the construction above gives the degrader for the flagged extension.
(ii) For Pauli/Clifford covariant noise that is degradable in a parameter regime, the same construction applies.
Appendix A Channel-capacity comparisons for AD/GADC and depolarizing channels
This section consolidates standard bounds used throughout quantum Shannon theory and places the numerical comparisons in Figs. 4, 5, and 6 into a compact theoretical frame. I state the precise theorems I invoke in our plots and include self-contained proofs or exact literature pointers.
A.1 Preliminaries
For a channel $\mathcal{N}_{A\to B}$, the coherent information is
$I_{c}(\mathcal{N},\rho_{A}):=H(\mathcal{N}(\rho_{A}))-H(\mathcal{N}^{c}(\rho_{A})),$ (A.1)
and the channel coherent information is $I_{c}(\mathcal{N}):=\max_{\rho_{A}}I_{c}(\mathcal{N},\rho_{A})$. The (unassisted) quantum capacity is
$Q(\mathcal{N})=\lim_{n\to\infty}\frac{1}{n}\,I_{c}\!\left(\mathcal{N}^{\otimes n}\right),$ (A.2)
while the entanglement-assisted classical capacity is the single-letter mutual information $C_{E}(\mathcal{N})=\max_{\rho_{RA}}I(R;B)_{\sigma}$, $\sigma_{RB}=(\mathrm{id}_{R}\!\otimes\!\mathcal{N})(\phi_{RA})$.
A.2 Results
Theorem A.1 (Quantum capacity of degradable channels).
If a channel $\mathcal{N}$ is degradable, i.e. there exists a CPTP map $\mathcal{D}$ such that $\mathcal{N}^{c}=\mathcal{D}\!\circ\!\mathcal{N}$, then the coherent information is additive $I_{c}(\mathcal{N}^{\otimes n})=n\,I_{c}(\mathcal{N})$ for all $n$, and hence
$Q(\mathcal{N})=I_{c}(\mathcal{N}).$ (A.3)
Proof.
The definition of degradability implies, for any state $\rho_{RA}$ and $n\!\geq\!1$, that the environment outputs of $\mathcal{N}^{\otimes n}$ are obtained from the channel outputs via a CPTP map $\mathcal{D}^{\otimes n}$. By data processing (monotonicity) of quantum relative entropy and strong subadditivity, the coherent information is superadditive and in fact additive on degradable channels; see, e.g., Theorems 13.6.1–13.6.2 in [21]. Therefore $I_{c}(\mathcal{N}^{\otimes n})=nI_{c}(\mathcal{N})$. Combining additivity with the regularized formula (A.2) (Devetak–Shor coding theorem and achievability via random stabilizer codes) yields $Q(\mathcal{N})=I_{c}(\mathcal{N})$. ∎
Proposition A.2 (Qubit amplitude damping (AD) is degradable iff $\gamma\!\leq\!\tfrac{1}{2}$).
Let $\mathcal{A}_{\gamma}$ be the qubit AD channel with Kraus operators $E_{0}=|0\rangle\!\langle 0|+\sqrt{1-\gamma}\,|1\rangle\!\langle 1|$, $E_{1}=\sqrt{\gamma}\,|0\rangle\!\langle 1|$. Then $\mathcal{A}_{\gamma}$ is degradable for $\gamma\in[0,\tfrac{1}{2}]$ and anti-degradable for $\gamma\in[\tfrac{1}{2},1]$. In particular, by Theorem A.1, $Q(\mathcal{A}_{\gamma})=I_{c}(\mathcal{A}_{\gamma})$ for $\gamma\leq\tfrac{1}{2}$, and $Q(\mathcal{A}_{\gamma})=0$ for $\gamma\geq\tfrac{1}{2}$.
Proof.
This is a special case of the qubit-channel degradability classification in [1]: amplitude-damping-type channels are degradable exactly for $\gamma\leq\tfrac{1}{2}$. An explicit degrading map can be constructed by expressing the complementary channel and solving for a CPTP map $\mathcal{D}$ such that $\mathcal{A}_{\gamma}^{c}=\mathcal{D}\!\circ\!\mathcal{A}_{\gamma}$; complete details and positivity constraints are given in [1, Sec. III]. Anti-degradability for $\gamma\geq\tfrac{1}{2}$ then implies zero capacity. ∎
Proposition A.3 (Reverse coherent information (RCI) bound).
For any channel $\mathcal{N}$, the reverse coherent information $I_{\mathrm{RCI}}(\mathcal{N}):=\max_{\rho_{RA}}[H(\rho_{R})-H((\mathrm{id}_{R}\!\otimes\!\mathcal{N})(\rho_{RA}))]$ is a generally-computable lower bound on $Q(\mathcal{N})$ and satisfies $I_{\mathrm{RCI}}(\mathcal{N})\leq I_{c}(\mathcal{N})$ for degradable $\mathcal{N}$.
Proof.
See [4] for the operational meaning of RCI and the proof that it is achievable (via reverse reconciliation protocols) and hence a lower bound to $Q$. For degradable channels, $I_{c}$ is single-letter and dominates RCI by data-processing. ∎
Proposition A.4 (Hashing bound for the qubit depolarizing channel).
For the Pauli depolarizing channel $\mathcal{D}_{p}(\rho)=(1-p)\rho+\tfrac{p}{3}(X\rho X+Y\rho Y+Z\rho Z)$, the hashing bound (one-shot coherent information) equals
$I_{c}(\mathcal{D}_{p})=1-H_{2}(p)-p\log_{2}3,$ (A.4)
yielding the standard achievable-rate curve used in Fig. 6.
Proof.
Because $\mathcal{D}_{p}$ is Pauli and unital, the optimizing input for $I_{c}$ can be taken as the maximally mixed state. A direct evaluation gives the spectrum of the output and its complement, leading to the closed form stated; see [21, Sec. 24.7]. ∎
Theorem A.5 (SDP / max-Rains upper bound).
Let $R_{\max}(\mathcal{N})$ denote the max-Rains information of a channel $\mathcal{N}$. Then the quantum capacity assisted by PPT-preserving codes is upper bounded as $Q^{\mathrm{PPT}}(\mathcal{N})\leq R_{\max}(\mathcal{N})$, which in turn upper bounds $Q(\mathcal{N})$. Moreover, $R_{\max}(\mathcal{N})$ is computable via semidefinite programming and is additive.
Proof.
This is established in [17, 20] using a meta-converse based on generalized divergences and the Rains bound on distillable entanglement. The additivity implies a single-letter efficiently-computable upper bound. ∎
Figure 4: Amplitude damping (AD): comparison of $I_{c}$, reverse coherent information (RCI), entanglement-assisted capacity $C_{E}$, and an SDP/max-Rains upper bound, as a function of damping parameter $\gamma$. The vertical line marks the anti-degradable threshold $\gamma=\tfrac{1}{2}$ (Prop. A.2).
Figure 5: Panel comparison for AD (left) and GADC at $N_{\mathrm{th}}\!\in\!\{0.0,0.1,0.2\}$ (right panels): $I_{c}$, RCI, $C_{E}$, and an SDP/max-Rains-style upper bound. The dashed line at $\gamma=\tfrac{1}{2}$ marks the AD anti-degradable threshold.
Figure 6: Qubit depolarizing channel: hashing lower bound $I_{c}(\mathcal{D}_{p})$ and an illustrative SDP/max-Rains upper bound.